### v1

@@ V Observational possibilities of the light curves
We have theoretically modeled a single pulse of energy that can be generated from magnetic reconnection in cosmic string wakes. We have then done a superposition of such pulses for random relativistic velocities of merger to obtain a synthetic light curve for a GRB. We would like to highlight that the burst duration and the energy released from the merging shocks depends primarily on the length of the magnetic reconnection region. This in turn depends on the magnetic field lengthscale. Since the magnetic field is generated in the wake of the cosmic string, the magnetic lengthscale would depend on the length of the wake. As mentioned previously, the magnetic fields are along the $z-$direction. They are however squeezed in the $y-$direction due to the flow of the plasma around the cosmic string. The velocity perturbation in the $y-$direction comes from the conical spacetime of the cosmic string. This space time has a deficit angle which is proportional to $G\mu$. Here, the $\mu$ or mass per unit length depends on the symmetry breaking scale which generated the cosmic string in the first place. We would now like to check whether any cosmic string model would indeed generate energies in the ranges where GRB’s have been observed.
@@
As we have mentioned before, we have considered the parameters of an Abelian Higgs cosmic string. We have calculated the reconnection length from the parameters of the wake of the cosmic string. For different values of $z$, this turns out to be different. Though our simulations are based on the energy density, one can relate the energy density to the fluence associated with a GRB. The fluence associated with these light curves can be found out by multiplying with the length associated with the wakes which are of the order of $10^{10}$ cm or less, so the fluences are found to be in the range of $\sim 10^{-5}$ erg cm-2 or less. For $z=30$, the timescale is found out to be $\Delta t=1.14$ sec. The collision time is therefore $\triangle t<2$ sec for different Lorentz factors corresponding to the two outgoing shock waves.
@@
We have studied the available catalogue of GRBs to compare our energy scales with observed bursts for different redshifts. The table below shows the various high-redshift GRBs observed for short duration GRB’s in the BATSE satellite. This is not a exhaustive list of short duration GRB’s but a token list to show the burst duration and fluence currently being observed by various telescopes. As is seen from the table, the energy scales and the timescales both indicate that magnetic reconnection in cosmic string wakes can be considered to be the progenitor for a GRB.
No. GRBs ID Redshift Burst Duration Fluence
(z) ($T_{90}$) (sec.) (×$\times 10^{-6}$ erg cm-2)
1. 090429B 9.4 0.53 $\pm$ 0.09 0.31 $\pm$ 0.03
2. 090423 8.2 1.13 $\pm$ 0.12 0.59 $\pm$ 0.04
3. 080913 6.7 1.04 $\pm$ 0.13 0.56 $\pm$ 0.06
4. 220521A 5.6 1.81 $\pm$ 0.07 0.98 $\pm$ 0.12
5. 240419A 5.17 0.48 $\pm$ 0.22 0.08 $\pm$ 0.02
6. 090205 4.65 1.55 $\pm$ 0.31 0.19 $\pm$ 0.03
@@
Most of our detailed calculations are done for the redshift value of $z=30$. Our motivation for this was the previously predicted signatures of cosmic string wakes in the 21 cm redshift surveys branden1 at $z=30$. Recently some studies have also been done for lower values of $z$ with the limitation of $z>15$ Oscar_2021 . For the GRB emission, the most important change due to the redshift would be the value of the background plasma density. We have found that changes in the plasma density only changes the maximum energy associalted with the light pulse. The mechanism for the generation of the GRB remains the same irrespective of the redshift. In a future work, we plan to look at actual experimental data to try and fit the model postulated in this paper at different redshifts to try and determine a signature of a cosmic string which will also support previous predicted signatures using density inhomogeneities in cosmic string wakes.
VI Summary and Conclusions
@@
To summarize, we have worked out a simplified model for the generation of a light pulse from the collision of magnetized shocks in the wake of a cosmic string at different redshifts. Magnetic field generated in cosmic string wakes are oppositely directed on both sides of the wakes. As the string moves through the plasma, the magnetic lines are pushed together and magnetic reconnection can occur. This reconnection results in the generation of a large amount of kinetic energy. Different models of magnetic reconnection gives rise to multiple shocks. We show that the collision of these shocks can result in the generation of a strong pulse of energy. Multiple shocks colliding may give rise to a GRB light curve. Using the various paramaters of shock waves due to magnetic reconnection in cosmic string wakes we have modeled the light pulse and the synthetic light curves. We find a generic pattern for the GRB light curve. It depends primarily on the density of the background plasma and the magnetic field. There is a weak dependence on the ratio of the densities of the colliding shocks, similarly there is a weak dependency on the velocities of the colliding shocks.
@@
There are some basic assumptions we have made for the current work. We have only considered the collision of shocks in one direction and have assumed that there is no correlation in the multiple shock mergers. This means that multiple collision will just enhance the overall energy of the outburst. A more detailed simulation would include various other factors. It is possible that after the initial collision between two shocks, the merged shock will coast for some time before colliding with another shock generated from a different point. This will require changes in the timescales and will therefore change the light curve considerably. Since we have found that the calculated fluence and the burst duration are well within the ambit of GRB’s observed by various experimental groups, we would like to improve on these simulations in a later work. Our primary objective in this work was to show conclusively that magnetic reconnection in the wakes of cosmic strings can indeed be the progenitor for a GRB.
@@
Acknowledgment: We acknowledge funding from the DST-SERB Power Grant no.SPG/2021/002228, D.K is supported by the Power Grant no. SPG/2021.002228, of the Government of India.

### v2

@@ VI Observational possibilities of the light curves
We have theoretically modeled a single pulse of energy that can be generated from magnetic reconnection in cosmic string wakes. We have then done a superposition of such pulses for random relativistic velocities of merger to obtain a synthetic light curve for a GRB. We would like to highlight that the burst duration and the energy released from the merging shocks depends primarily on the length of the magnetic reconnection region. This in turn depends on the magnetic field lengthscale. Since, the magnetic field is generated in the wake of the cosmic string, the magnetic lengthscale would depend on the length of the wake. As mentioned previously, the magnetic fields are along the $Z$-direction. In our previous work Dilip_2023 , we have mentioned that cosmic string wakes are essentially planar wakes and so we are looking at a quasi-two dimensional process. The two dimensions considered here are the XYX-Y plane. The hot electron current is in opposite directions in the two dimensional plane, so basically for the quasi-two dimensional plane we have approximated $\vec{B}~=~-\hat{i}\frac{\partial N_{e}}{\partial z}\frac{\partial T}{\partial y}$ to map it on the two dimensional plane [For detail study check ref.Dilip_2023 ]. The magnetic lines are squeezed in the $Y$-direction due to the flow of the plasma around the cosmic string. The velocity perturbation in the $Y$-direction comes from the conical spacetime of the cosmic string. This space time has a deficit angle which is proportional to $G\mu$. Here, the $\mu$ is the mass per unit length of the string depends on the symmetry breaking scale which generated the cosmic string in the first place. We would now like to check whether any cosmic string model would indeed generate energies in the ranges where GRB’s have been observed.
@@
As we have mentioned before, we have considered the parameters of an Abelian Higgs cosmic string. All the constraints on the cosmic string parameters are based on the results from Planck Planck_2016 . We have calculated the reconnection length from the parameters of the wake of the cosmic string. For different values of $z$, this turns out to be different. Though our simulations are based on the energy density, one can relate the energy density to the fluence associated with a GRB. The fluence associated with these light curves can be found out by multiplying with the length associated with the wakes which are of the order of $10^{10}$ cm or less, so the fluences are found to be in the range of $\sim 10^{-5}$ erg cm-2 or less. For redshift $z=6.7$, the timescale is found out to be $\Delta t=1.61$ sec. The collision time is therefore $\Delta t<2$ sec for different Lorentz factors corresponding to the two outgoing shock waves. Thus, the obtained result may be for short-duration GRBs. One of the important factors in GRB’s is the electromagnetic emission band predicted by the model. The crucial part of our work is the energy generated in the reconnection region as that is the energy which produces the shocks. As long as this energy is high enough to produce a GRB, the shocks emanating from these points will have this energy as a minimum. So we calculate the total energy generated in the string wake due to the magnetic reconnection. We base our calculation on the Petschek model but emphasis that there are other models which can generate even faster reconnections. The width of the reconnection region ($\delta$) is given by Petschek
$\delta=\frac{c^{2}}{4\pi\sigma V_{A}M_{0}},$ (13)
where, $M_{0}$ is calculated from the diffusion region as,
$M_{0}^{2}=\frac{c^{2}}{8\pi\sigma V_{A}y^{*}},$ (14)
where, $y^{*}$ is the length of reconnection region, $V_{A}=\frac{B}{\sqrt{4\pi\rho}}$. We consider values at $z=6.7$. Here we have taken $B$ to be the equipartition magnetic field at the given redshift $\sim 1.95\times 10^{-2}~eV^{2}$, $\rho$ is plasma density at that redshift $\sim 10^{-10}~eV^{4}$ and $\sigma$ is the conductivity of the plasma $\sim 0.33~cm^{-1}$ which has been obtained before. Thus, $y^{*}=l$ is length of reconnection region obtained is $\sim 1.40\times 10^{10}~cm$ from the cosmic string wake parameters. Hence, we obtain, $M_{0}\simeq 10^{-6}$ and substituting this in the width of the reconnection region, we obtain, $\delta\sim 10^{2}~cm$.
So we have obtained the energy density from magnetic reconnection in cosmic string wakes as, $\mathcal{E}\sim 10^{-15}~ergcm^{-3}$. Since, $1~erg=6.24\times 10^{2}~GeV$. Thus, $\mathcal{E}\sim 6.24\times 10^{-4}~eVcm^{-3}$. Now, for the total energy emitted from the reconnection region geometry, we use, $E=\mathcal{E}\times V$, where, $V$ is volume of the reconnection region. Thus, we obtain the total energy as $E\sim 10^{2}MeV$. This is the minimum energy band. As mentioned previously, magnetic fields can get amplified in cosmic string wakes which means that the actual magnetic field in the calculation should be at least $10$ times higher than the magnetic field we have considered. This will enhance the total energy generated in the magnetic reconnection region. This will in turn increase the energy of the shocks and therefore the total energy emitted.
@@
We have studied the available catalogue of GRBs to compare our energy scales with observed bursts for different redshifts. The table below shows the various high-redshift GRBs observed for short duration GRB’s. This is not a exhaustive list of short duration GRB’s but a token list to show the fluence and rest-frame burst duration ($T_{90}$) currently being observed by various satellites like FERMI, SWIFT mentioned in individual references for each GRBs. The rest-frame duration is calculated as the observed burst duration divided by $(1+z)$, accounting for cosmological time dilation. As we have also done the calculation for cosmic string wake in the rest frame. This correction allows for a fair comparison of GRB durations across different redshifts by removing the stretching effect due to the expansion of the universe. As is seen from the table, the energy scales and the timescales both indicate that magnetic reconnection in cosmic string wakes can be considered to be the progenitor for a GRB.
S.No. GRBs ID Redshift Burst Duration ($T_{90}$) Observed Fluence Ref.
(sec.) ($\times 10^{-6}$ erg cm-2)
1. 090429B 9.4 0.53 $\pm$ 0.09 0.31 $\pm$ 0.03 Cucchiara_2011
2. 090423 8.1 1.13 $\pm$ 0.12 0.59 $\pm$ 0.04 Salvaterra_2009
3. 080913 6.7 1.04 $\pm$ 0.13 0.56 $\pm$ 0.06 Greiner_2009
4. 220521A 5.6 1.81 $\pm$ 0.07 0.98 $\pm$ 0.12 Minaev_2022
5. 240419A 5.17 0.48 $\pm$ 0.22 0.08 $\pm$ 0.02 Parsotan_2024
6. 090205 4.65 1.55 $\pm$ 0.31 0.19 $\pm$ 0.03 DAvanzo_2010
We have also looked at the published data and made an attempt to fit the data from our model. The model does not incorporate many of the finer points of the plasma dynamics that occur in a cosmic string wake but even then it can fit the data to a great extent. We have calculated the flux and matched it with the data from the reference chang . This is shown in Fig.9. We have fitted the data from the GRB 081215. As is seen in the figure, we do not get a perfect match. For data fitting, we have mostly varied the shock collision velocities and the magnetic field as these are the unknown parameters of our model. An interesting observation was that a constant magnetic field does not fit the data well. This is not surprising as the magnetic field in the cosmic string wake will not necessarily be a constant field. The collision timescale and length of the reconnection region remain fixed by the parameters of the cosmic string. But due to diffusion processes, the magnetic field does not remain the same as the shocks travel in an outward direction. We have varied the free parameters $\alpha$ and $\beta$ to fit the data from the GRB. This fit is shown in fig. 9. The fit that we have obtained is for a value of $\alpha=13$ and $\beta=2$. Thus the fit strongly depends on the way the magnetic field decays. We have only fitted the dominant pulse. A best fit for all the data will require including other dissipative forces.
Figure 9: The graph shows the best fit for the data from GRB 081215 to our model. The $\alpha$ parameter is quite high with the current fit being given for $\alpha=13$ and $\beta=2$.
@@
Most of our detailed calculations are done for the redshift value of $z=6.7$. Recently some studies on cosmic strings have been done for lower values of $z$ with the limitation of $z>15$ Oscar_2021 . For the GRB emission, the most important change due to the redshift would be the value of the background plasma density. We have found that changes in the plasma density only changes the maximum energy associated with the light pulse. The mechanism for the generation of the GRB remains the same irrespective of the redshift. In a future work, we plan to improve the model by incorporating other features of the plasma. This will enable us to fit the experimental data with greater accuracy. We believe that in addition to the other signatures of a cosmic string, this will help us in determining the presence of a cosmic string in the current universe.
VII Summary and Conclusions
To summarize, we have worked out a model for the generation of a light pulse from the collision of shocks in the magnetized wake of a cosmic string at different redshifts. Gamma-ray bursts (GRBs) arising from cosmic strings have been studied in the literature in the context of superconducting cosmic strings before. However in our model we consider non-superconducting cosmic strings where the magnetic fields arise due to the Biermann battery mechanism. This will lead to one significant difference. The GRB’s generated from the cusps of super conducting strings are strongly collimated in a narrow region. The GRB’s generated in the wake region will not be so constrained. Though the energy would appear to be collimated but the opening angle would be larger than in the case of the superconducting cosmic strings. In the current work, we have not calculated the event rate of the GRB’s from the cosmic string wake. This is because unlike the superconducting cosmic string case, there are many more parameters that need to be evaluated precisely before an event rate can be calculated. Primary amongst this is the understanding of the exact kind of magnetic reconnection that is happening in the cosmic string wake. This study itself will require large scale MHD simulations which are currently beyond the scope of this work. We plan to follow it up later and we will try to include all the possible parameters in the collision model. There are other ways in which we can also pinpoint whether the GRB is from a cosmic string wake. As mentioned in the Introduction section, the wakes will leave an imprint on the 21cm redshift survey maps. Mapping these imprints with GRB bursts may lead us to a cosmic string wake. Since any cosmic string has a wake structure behind it, we do not need to look at Superconducting strings only. This we believe increases the chances of identifing cosmic strings from the current observational data. We will now conclude with a brief summary of the work.
@@ VII Summary and Conclusions
Magnetic field generated in cosmic string wakes are oppositely directed on both sides of the wakes. As the string moves through the plasma, the magnetic field lines are pushed together and magnetic reconnection can occur. This reconnection results in the generation of a large amount of kinetic energy. Different models of magnetic reconnection gives rise to multiple shocks. We show that the collision of these shocks can result in the generation of a strong pulse of energy. Multiple shocks colliding may give rise to a GRB light curve. Using the various parameters of shock waves due to magnetic reconnection in cosmic string wakes we have modeled the light pulse and the synthetic light curves. We find a generic pattern for the GRB light curve. It depends primarily on the density of the background plasma and the magnetic field. There is a weak dependence on the ratio of the densities of the colliding shocks, similarly there is a weak dependency on the velocities of the colliding shocks. We have fitted the data of an observed GRB to our model. Though the fit is not perfect, it shows that the generic pattern of the GRB is reflected in the data too. This indicates that the model can be improved to fit the data from existing GRB’s by incorporating further details of the plasma dynamics.
@@
Since this is the first work on the GRB’s from non-superconducting cosmic string we have considered some basic assumptions for the current work. We have only considered the collision of shocks in one direction and have assumed that there is no correlation in the multiple shock mergers. This means that multiple collision will just enhance the overall energy of the outburst. A more detailed simulation would include various other factors. It is possible that after the initial collision between two shocks, the merged shock will coast for some time before colliding with another shock generated from a different point. This will require changes in the timescales and will therefore change the light curve considerably. As we have mentioned in the text, the cosmic string wake is bounded by strong shocks on both sides. This will lead to cases of shock reflections and the collision with a reflected shock. The shocks are also oblique in nature so it may not be a smooth merger. We have not considered any of these cases. We have not considered several other cases such as Mach stems, all of which can occur during shock collisions. We believe that detailed studies are necessary to improve the lightcurve further. Since we have found that the calculated fluence and the burst duration are well within the ambit of GRB’s observed by various experimental groups, we would like to incorporate some more features of shock collisions in these simulations in a later work. Our primary objective in this work was to show conclusively that magnetic reconnection in the wakes of cosmic strings can indeed be the progenitor for a GRB.
@@
Acknowledgment: We acknowledge funding from the DST-SERB/ANRF Power Grant no.SPG/2021/002228, D.K is supported by the Power Grant no. SPG/2021.002228, of the Government of India.